# Reasoning to Edit: Hypothetical Instruction-Based Image Editing with Visual Reasoning

### v1

@@ C More Qualitative Results
To further demonstrate the editing performance of ReasonBrain compared to SOTA methods, we present additional qualitative results in Fig. A2 and Fig. A3. It is evident that ReasonBrain consistently outperforms other SOTA approaches, producing more coherent and visually plausible edits. These results highlight ReasonBrain’s ability to reason effectively from hypothetical instructions and generate outputs that closely align with the transformations implied by the underlying reasoning.
D Broader Impact and Ethics Statement
@@
![[Uncaptioned image]](extracted/6590845/figs/emjio_right.png)
![[Uncaptioned image]](extracted/6590845/figs/emojio_x.png)
![[Uncaptioned image]](extracted/6590845/figs/ReasonP2P1.png)
![[Uncaptioned image]](extracted/6590845/figs/ReasonP2P2.png)
![[Uncaptioned image]](extracted/6590845/figs/editworld_source.png)
![[Uncaptioned image]](extracted/6590845/figs/editworld_target.png)
![[Uncaptioned image]](extracted/6590845/figs/smartedit_source.png)
![[Uncaptioned image]](extracted/6590845/figs/smartedit_target.png)
![[Uncaptioned image]](extracted/6590845/figs/reason_source.png)
![[Uncaptioned image]](extracted/6590845/figs/reason_target.png)

### v2

@@ H More Qualitative Results
To further demonstrate the editing performance of ReasonBrain compared to SOTA methods, we present additional qualitative results in Fig. A2 and Fig. A3, Fig. A4 and Fig. A5. It is evident that ReasonBrain consistently outperforms other SOTA approaches, producing more coherent and visually plausible edits. These results highlight ReasonBrain’s ability to reason effectively from hypothetical instructions and generate outputs that closely align with the transformations implied by the underlying reasoning.
I Broader Impact and Ethics Statement
@@
![[Uncaptioned image]](figs/emjio_right.png)
![[Uncaptioned image]](figs/emojio_x.png)
![[Uncaptioned image]](figs/ReasonP2P1.png)
![[Uncaptioned image]](figs/ReasonP2P2.png)
![[Uncaptioned image]](figs/editworld_source.png)
![[Uncaptioned image]](figs/editworld_target.png)
![[Uncaptioned image]](figs/smartedit_source.png)
![[Uncaptioned image]](figs/smartedit_target.png)
![[Uncaptioned image]](figs/reason_source.png)
![[Uncaptioned image]](figs/reason_target.png)